OpenStudy (anonymous):

how do I work out (5/2)/2

OpenStudy (king):

\[\frac{5}{2}\div 2\] =\[\frac{5}{2}\times \frac{1}{2}\] =\[\frac{5}{4}\]

OpenStudy (anonymous):

5/4